OpenStudy (anonymous):

what is 14 base 7 to base 10

OpenStudy (zappy620):

This has to do with logarithms correct?

OpenStudy (bigdaddyharsha):

11.8313725602

Directrix (directrix):

14 base 7 is 1*7 + 4*1 = 7 + 4 = 11 base 10 ----------- 11.8313725602 is NOT correct @varunvarsity